OpenStudy (cookie_monster):

If a sample of 26 marbles has a mass of 55 grams and a volume of 22 milliliters, what is the density of that sample?

OpenStudy (cookie_monster):

I got 2.5 but I don't know if it's actually correct.

OpenStudy (anonymous):

i think it is 33. 55-22 is 33. i think thats right

OpenStudy (gudden):

density = mass over volume so u are correct :D

OpenStudy (bibby):

\[d = \frac{ m }{ v }\] where m = 55g and v = 22ml \[\large = \frac{ 55g }{ 22ml } = 2.5\dfrac{ g }{ ml }\]

OpenStudy (cookie_monster):

Thank you peoples!!

OpenStudy (gudden):

ur welcome...:D